Two numbers (x & y) differ by 4, but their product is 96. What are the two numbers?
I tried:
x(y+4)=96
xy+4x-96=0
(x+12)(y-8)
x=-12 y=8
It's not an age question but I thought it might be able to be done in a similar way.

Answer by Edwin McCravy(20055)


Two numbers (x & y) differ by 4,
x-y = 4     <--first equation

but their product is 96.
xy = 96     <--second equation

Solve the first equation for x

x = y+4

Substitute in the second equation:

(y+4)y = 96

y(y+4) = 96

y²+4y-96 = 0

(y+12)(y-8) = 0

y+12=0;  y-8=0
   y=-12   y=8

if y = -12, we substitute in

x = y+4      
x = -12+4
x = -8

So one answer is x = -8 and y = -12

if y = 8, we substitute in

x = y+4       
x = 8+4
x = 12

So the other answer is x = 12 and y = 8

Edwin